# Vision-based 3D occupancy prediction in autonomous driving: a review and outlook

### v1

@@ 1 Introduction
3D object detection, as a fundamental component of autonomous driving perception systems, has drawn widespread attention from researchers, resulting in three branches based on diverse sensor data modalities: image-based methods [1, 2, 3, 4, 5, 6], LiDAR-based methods [7, 8, 9, 10, 11, 12], and multi-modal methods [13, 14, 15, 16, 17, 18]. While LiDAR-based and multi-modal methods can achieve relatively strong detection performance, image-based methods are favored for practical deployment due to their lower economic costs and better real-time capabilities. Recently, the emergence of the Bird’s-Eye-Vie (BEV) representation paradigm [19, 20, 21, 22, 23] has improved the accuracy of vision-centric 3D object detection, further narrowing the performance gap.
@@
In this paper, we comprehensively review the vision-based 3D occupancy prediction methods for autonomous driving applications and systematically categorize existing methods into three groups: feature enhancement methods, deployment-friendly methods and label-efficient methods. For feature enhancement, we further summarize the feature representation methods based on Bird’s Eye View (BEV), Tri-Perspective View (TPV) and 3D Voxel. From computational friendliness, we conduct an in-depth analysis from two viewpoint: perspective decomposition and coarse-to-fine paradigm. As for label efficiency, we methodically categorize relevant research into annotation-free methods and LiDAR-free methods. Finally, we provide some inspiring future outlooks for 3D occupancy prediction from the perspectives of data, methodology, and task.
@@
  • This paper structurally summarizes vision-based 3D occupancy prediction methods from three views: feature-enhanced, computation-friendly and label-efficient approaches, and provide in-depth analysis and comparison for different categories of methods.
@@ 2.3 Datasets
Table 1: Comparison of datasets for 3D occupancy prediction. Mult-view=$\checkmark$ represents datasets that use mult-view inputs. $C$, $D$, $L$ denote camera, depth and LiDAR.
@@ 3.2 TPV-based methods
While BEV-based representations have certain advantages compared to images, as they essentially provide a top-down projection of 3D space, they inherently lack the ability to describe the fine-grained 3D structure of scenes using only a single plane. Tri-Perspective View (TPV) based methods utilize three orthogonal projection planes to model the 3D environment, further enhancing the representation capability of visual features for occupancy prediction. First, image features are extracted from visual inputs using a 2D backbone network. Subsequently, these image features are elevated to a three-view space, and ultimately, 3D occupancy prediction is accomplished based on the feature representation from the three projection viewpoints. The illustration of BEV-based methods is shown in Figure 7.
@@
Apart from BEV features, TPVFormer [40] also generates the features in the front view and side view in the same way. Each plane models the 3D environment from a different viewpoint, and their combination provides a comprehensive description of the whole 3D structure. Specifically, to obtain the feature of a point in 3D space, we first project it onto each of the three planes and use bilinear interpolation to obtain the feature of each projected point. Then, we sum up the three projected features as the composite feature of the 3D point. Thus, TPV representation can describe the 3D scene at arbitrary resolutions and generate different features for different points in 3D space. It further propose a Transformer-based encoder (TPVFormer) to effectively obtain TPV features from 2D images and perform image cross-attention between the TPV grid queries and the corresponding 2D image features to elevate 2D information to 3D space. Finally the cross-view mixing attention among TPV features achieve interaction between the three planes. The overall architecture of TPVFormer is shown in Figure 8.
@@
Considering TPVFormer only focuses on spatial cues while neglecting temporal cues, S2TPVFormer [68] further extends the temporal fusion ability of TPVFormer. It adopts TPV as a latent ego spatial representation and proposes Temporal Cross-View Hybrid Attention, which first interacts with the current BEV features with previous BEV features. After that, the front and side view features will interact with the updated BEV features to fusion temporal information. Taking into account the non-uniform distribution of LiDAR point clouds in space, with denser points nearby and sparser points further away, PointOcc [69] propose establishing a three-view representation of point clouds in cylindrical coordinates to achieve finer-grained modeling of nearby regions, and employing spatial grouping pooling during the projection process to better preserve 3D structural information. Specifically, given the input point cloud, it is first transformed into tri-view features through voxelization and spatial pooling operations in cylindrical coordinates. Then, a 2D image backbone is employed to encode each tri-view feature plane, and multi-scale features from each plane are aggregated to decode high-resolution tri-view feature representations. For any query point in 3D space, it is projected onto each tri-view plane through cylindrical coordinate transformation to sample the corresponding features. These features from the tri-view planes are then aggregated using summation to form the final 3D feature representation.

### v2

@@ 1 Introduction
3D object detection, as a fundamental component of autonomous driving perception systems, has drawn widespread attention from researchers, resulting in three branches based on diverse sensor data modalities: image-based methods [1, 2, 3, 4, 5, 6], LiDAR-based methods [7, 8, 9, 10, 11, 12], and multi-modal methods [13, 14, 15, 16, 17, 18]. While LiDAR-based and multi-modal methods can achieve relatively strong detection performance, image-based methods are favored for practical deployment due to their lower economic costs and better real-time capabilities. Recently, the emergence of the Bird’s-Eye-View (BEV) representation paradigm [19, 20, 21, 22, 23] has improved the accuracy of vision-centric 3D object detection, further narrowing the performance gap.
@@
In this paper, we comprehensively review the vision-based 3D occupancy prediction methods for autonomous driving applications and systematically categorize existing methods into three groups: feature enhancement methods, deployment-friendly methods and label-efficient methods. For feature enhancement, we further summarize the feature representation methods based on Bird’s Eye View (BEV), Tri-Perspective View (TPV) and 3D Voxel. From computational friendliness, we conduct an in-depth analysis from two viewpoints: perspective decomposition and coarse-to-fine paradigm. As for label efficiency, we methodically categorize relevant research into annotation-free methods and LiDAR-free methods. Finally, we provide some inspiring future outlooks for 3D occupancy prediction from the perspectives of data, methodology, and task.
@@
  • This paper structurally summarizes vision-based 3D occupancy prediction methods from three views: feature-enhanced, computation-friendly and label-efficient approaches, and provides in-depth analysis and comparison for different categories of methods.
@@ 2.3 Datasets
Table 1: Comparison of datasets for 3D occupancy prediction. Mult-view=$\checkmark$ represents datasets that use mult-view inputs. $C$, $D$, $L$ denote camera, depth and LiDAR. GO is short for General Object.
@@ 3.2 TPV-based methods
While BEV-based representations have certain advantages compared to images, as they essentially provide a top-down projection of 3D space, they inherently lack the ability to describe the fine-grained 3D structure of scenes using only a single plane. Tri-Perspective View (TPV) based methods utilize three orthogonal projection planes to model the 3D environment, further enhancing the representation capability of visual features for occupancy prediction. First, image features are extracted from visual inputs using a 2D backbone network. Subsequently, these image features are elevated to a three-view space, and ultimately, 3D occupancy prediction is accomplished based on the feature representation from the three projection viewpoints. The illustration of TPV-based methods is shown in Figure 7.
@@
Apart from BEV features, TPVFormer [40] also generates the features in the front view and side view in the same way. The triple views are orthogonal to each other and squeeze the 3D space information along different axes. The whole 3D structure of the driving scene can be recovered by combining the TPV features. To retrieve the feature of a certain position in the 3D space from TPV features, the 3D coordinates are first projected on each of the views. After that, bilinear sampling is utilized to obtain the features from the triple views, which are summarized as the composite features. This mechanism makes 2D TPV features able to flexibly recover the 3D space without losing too much information. In practice, the TPV features are generated by utilizing a Transformer-based encoder, which consists of several cross-attention between TPV grid queries and 2D image features. Finally, the cross-view hybrid attention among TPV features achieves interaction between the three planes, improving the capability of the TPV feature to represent the whole scene. The overall architecture of TPVFormer is shown in Figure 8.
@@
Considering TPVFormer only focuses on spatial cues while neglecting temporal cues, S2TPVFormer [68] further extends the temporal fusion ability of TPVFormer. It adopts TPV as a latent ego spatial representation and proposes Temporal Cross-View Hybrid Attention, which first interacts with the current BEV features with previous BEV features. After that, the front and side view features will interact with the updated BEV features to fusion temporal information. Taking into account the non-uniform distribution of LiDAR point clouds in space, with denser points nearby and sparser points further away, PointOcc [69] proposes establishing a three-view representation of point clouds in cylindrical coordinates to achieve finer-grained modeling of nearby regions, and employing spatial grouping pooling during the projection process to better preserve 3D structural information. Specifically, given the input point cloud, it is first transformed into tri-view features through voxelization and spatial pooling operations in cylindrical coordinates. Then, a 2D image backbone is employed to encode each tri-view feature plane, and multi-scale features from each plane are aggregated to decode high-resolution tri-view feature representations. For any query point in 3D space, it is projected onto each tri-view plane through cylindrical coordinate transformation to sample the corresponding features. These features from the tri-view planes are then aggregated using summation to form the final 3D feature representation.